OpenStudy (erinweeks):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer.

OpenStudy (erinweeks):

\[\log_{6} x ^{2} = \log_{6} (5x + 36)\]

OpenStudy (erinweeks):

A. {3/2} B. {9} C. Ø D. {9, -4}

OpenStudy (erinweeks):

@jim_thompson5910 can you help?

OpenStudy (goformit100):

Madam, can I help you ?

OpenStudy (erinweeks):

yes you may ((: @goformit100

OpenStudy (goformit100):

use the basic formula of log.

OpenStudy (erinweeks):

whats the formula?

OpenStudy (erinweeks):

and i really dont understand how to use the formulas

OpenStudy (mandre):

log6x2=log6(5x+36) Therefore x^2 = 5x + 36 Solve for x

OpenStudy (erinweeks):

are both sides suppose to be the same

OpenStudy (mandre):

If you have Logx Y = logx Z then Y = Z bring all terms to one side then solve it.

OpenStudy (erinweeks):

im confused i dont know how to bring it over

OpenStudy (mandre):

Subtract 5x both sides then do the same for 36. If you have it in the form ax2 + bx + c = 0 then you can factorize.

OpenStudy (erinweeks):

you cant subtract 5x from the left side .. could you show me how to do it so i understand

OpenStudy (mandre):

x^2 = 5x + 36 x^2 - 5x = 5x + 36 - 5x x^2 - 5x = 36

OpenStudy (erinweeks):

okay so now i have make the left side equal 36?

OpenStudy (mandre):

Nooo Do the same with 36 that you did with 5x.

OpenStudy (erinweeks):

so than it'd be x^2 - 5x = 36 - 36 x^2 - 5x - 36 = ?

OpenStudy (mandre):

You tell me :)

OpenStudy (erinweeks):

9!

OpenStudy (erinweeks):

?

OpenStudy (mandre):

What is 9?

OpenStudy (erinweeks):

the answer...

OpenStudy (mandre):

How did you get 9?

OpenStudy (erinweeks):

well 9 makes the equation 0 .. that not what i was going for?

OpenStudy (mandre):

Do you know how to factorize? That is what you must do. You need to work out the answer, not guess.

OpenStudy (erinweeks):

i did i did 9^2 = 81 than 5*9=45 81 - 45 - 36 = 0

OpenStudy (mandre):

You need to show how you got the 9. What if it was not a multiple choice question?

OpenStudy (erinweeks):

your right sorry .. but did i do it correct?

OpenStudy (mandre):

You need to factorise x^2 - 5x - 36 = 0

OpenStudy (erinweeks):

i though i just did....

OpenStudy (erinweeks):

5x^2 - 36 = 0 is that right?

OpenStudy (mandre):

Nope. If you did you would know that 9 is not the only answer.

OpenStudy (erinweeks):

did i just factorize it above?

OpenStudy (mandre):

No.

OpenStudy (erinweeks):

how do i factorize

OpenStudy (mandre):

Example: If you have x^2 - 4x -12 = 0 If you factorize you get (x - 6)(x + 2) = 0

OpenStudy (erinweeks):

so (x-9)(x+4) = 0

OpenStudy (mandre):

There you go.

OpenStudy (erinweeks):

okay so the answer would be D than

OpenStudy (erinweeks):

right

OpenStudy (mandre):

Yes.